A solvent passes through a chromatography column in 2.85 min, but the solute requires 9.43 min.

What is the retention factor, k?k=

What fraction of the time does the solute spend in the mobile phase in the column?

tfraction in mobile phase=

The volume of the stationary phase is 0.115 times the volume of the mobile phase in the column (Vs=0.115Vm).

What is the partition coefficient, K, for this system?

K=
Answer #1

Retention factor is the distance moved by solute decided by distance moved by solvent.

Solvent takes 2.85 min

Solute takes 9.43 min

Retention factor is 9.45-2.85÷2.85=2.31

Fraction time in mobile phase =9.43÷2.85=3.308

Partition coificiant is concentration of solute in stationary state÷ concentration of solute in mobile state.

Wich is equal to volume in mobile state ÷ volume in stationary state.=1÷0.115=8.695